# Step1X-Edit: A Practical Framework for General Image Editing

### v2

@@ 1 Introduction
Image editing with natural language instructions has become an increasingly important task in vision-language research. It offers intuitive interaction for end users while posing unique technical challenges: understanding nuanced semantics, precisely localizing regions to edit, and preserving image fidelity. While diffusion models [37, 5, 10, 36] have dramatically improved image generation quality, the existing design by integrating text encoder, e.g., CLIP [39] and T5 [40], with diffusion transformer often struggles with following editing instruction to maintain alignment between input image and edit instruction, especially when edit instructions are subtle or compositional.
@@
Building on this dataset, we propose, Step1X-Edit, a unified image editing model that combines the strong semantic reasoning of Multimedia Large Language Model (MLLM), e.g., Qwen-VL [3], with a DiT-style diffusion architecture. The reference image and editing prompts can be processed by the MLLM to generate a target image latent condition which will be integrated with the diffusion model to obtain the output image. Our approach maintains a good balance between reference image reconstruction and editing prompt following. To train the model, we start from a text-to-image model to retain aesthetic quality and visual consistency, which can be easily replace by the existing text-to-image models like SD3 [10] and FLUX[5]. To evaluate the existing editing models, we introduce a new benchmark named GEdit-Bench. By carefully collecting the images and editing prompts, GEdit-Bench ensures both real-world editing requirements and the diversity of the editing prompts. The experiments on GEdit-Bench validate that Step1X-Edit outperforms existing open-source baselines with a large-margin and approaches the performance of leading proprietary models, e.g., GPT-4o.
@@ 2.1 Controllable Image Generation and Edit
Diffusion models have become the dominant approach for high-fidelity image synthesis, offering strong capabilities in photorealism, structural consistency, and diversity. Beginning with DDPM [17] and DDIM [46], and further advanced by Latent Diffusion [42, 37], diffusion models operate in latent spaces for improved scalability. With the introduction of DiT architectures [36], diffusion models have made significant strides in generalization, image quality, and knowledge capacity, becoming the predominant architecture in modern image generation [1, 5]. Based on the above text2image models, ControlNet [63], T2I-Adapter [32], and OmniControl [57] inject spatial or task-specific control into the generation process. BrushNet [21], PowerPaint [68], and FLUX-Fill [6] further improve inpainting quality and versatility. Despite these advances, diffusion models often rely on static prompts or fixed conditions and lack the capacity for multi-turn reasoning or flexible language alignment, limiting their application in open-ended editing scenarios.
@@
2.2 Unified Instruction-based Image Editing Models
Unified instruction-based image editing models aim to bridge semantic instruction understanding and precise visual manipulation. Early approaches such as Prompt-to-Prompt [16], InstructEdit [50], InstructPix2Pix [7], MagicBrush [62], and BrushEdit [25] adopt modular pipelines where MLLMs generate prompts, spatial cues, or synthetic instruction-image pairs to guide diffusion-based editing.
Recent works move toward tighter integration between instruction and generation. SmartEdit [19], X2I [30], RPG [58], AnyEdit [60], and UltraEdit [66] enhance multimodal interaction and instruction fidelity through improved model architectures, task-aware routing, and fine-grained editing capabilities. Meanwhile, unified generation and editing frameworks such as OmniGen [56], ACE [15], ACE++[31], and Lumina-OmniLV[38] consolidate diverse visual tasks under a single architecture. Methods like OminiControl [48], Qwen2VL-Flux [28], and MetaQueries [34] further explore efficient control integration and direct latent-level fusion between MLLMs and diffusion decoders.
@@ 3.2 Our Method
The extracted embeddings are then fed into a lightweight connector module, such as the token refiner [29, 22]. This module restructures the embeddings into a more compact textual feature representation, which is subsequently used as the text embedding input to the downstream DiT network. Furthermore, we calculate the mean of valid embeddings from Qwen. This mean value is then projected through a linear layer, generating a global visual guidance vector. By doing so, the image editing network can leverage Qwen’s enhanced semantic comprehension capabilities, enabling more accurate and context-aware editing operations.
To effectively train the Token Refiner and enable rich cross-modal conditioning, we carefully design our feature aggregation strategy. Compared to approaches such as FLUX-Fill [6], which uses channel concatenation, and methods like SeedEdit [45], which introduce additional causal self-attention mechanisms, we follow OminiControl [48] and adopt token concatenation to better balance responsiveness to editing instructions with the preservation of fine-grained image details. During training, the reference image is encoded by a VAE encoder, and its latent features are linearly projected into reference image tokens. After processing, the image tokens are concatenated with noise image tokens along the token length dimension to construct the final visual input.
Recent research has also explored alternative approaches to enhance cross-model understanding by leveraging MLLMs. For instance, Qwen2VL-FLUX [28] introduces a pioneering approach by replacing the traditional T5 [40] text encoder in DiT-based text-to-image models with MLLMs to enhance multi-modal understanding and generation. However, it still retains T5 [40] for text encoding, which limits its ability to perform comprehensive cross-modal reasoning. In contrast, approaches like DreamEngine [8] leverages Qwen to align image and text modalities, with their features serving as external conditional inputs for SD3.5 [1] in the image synthesis. This method establishes a shared representation space that facilitates a more coherent process of generation and understanding. Nevertheless, in DreamEngine, the challenge remains in fully capturing the fine-grained details of reference images using only MLLM features.
@@ 4.1 GEdit(Genuine Edit)-Bench
Table 2: Quantitative evaluation on GEdit-Bench-EN. All metrics are reported as higher-is-better ($\uparrow$). The Intersection subset reflects the subset of prompts where all methods return valid responses with a total of 434 instances; the Full set includes all the 606 instances. G_SC, G_PQ, and G_O refer to the metrics evaluated by GPT-4.1, while Q_SC, Q_PQ, and Q_O refer to the metrics evaluated by Qwen2.5-VL-75B.
@@
Table 3: Quantitative evaluation on GEdit-Bench-CN. All metrics are reported as higher-is-better ($\uparrow$). The Intersection subset reflects the subset of prompts where all methods return valid responses with a total of 422 instances; the Full set includes all the 606 instances. G_SC, G_PQ, and G_O refer to the metrics evaluated by GPT-4.1, while Q_SC, Q_PQ, and Q_O refer to the metrics evaluated by Qwen2.5-VL-75B.
@@ 4.2.1 Evaluation on GEdit-Bench
Based on the GEdit-Bench, we evaluated a diverse range of image editing algorithms, covering state-of-the-art open-source solutions such as Instruct-Pix2Pix [7], MagicBrush [62], AnyEdit [60], and OmniGen [56], as well as proprietary algorithms like GPT-4o [33] [^1] The results are obtained based on ChatGPT APP in April 2025., doubao [45] [^2] The results are obtained based on doubao APP in April 2025., and Gemini2 Flash [14] [^3] The results are obtained in April 2025.. Following VIEScore [23], we adopt three metrics: SQ (Semantic Consistency), PQ (Perceptual Quality), and O (Overall Score). SQ assesses the degree to which the edited results conform to the given editing instruction, with a score ranging from 00 to $10$. PQ evaluates the naturalness of the image and the presence of artifacts, also using a scoring scale that ranges from 00 to $10$. The overall score is calculated based on these evaluations. To perform the automatic evaluation for VIEScore, we adopt the state-of-art MLLM model GPT-4.1444API access as of April 2025. Also, the evaluation based on the open-source model Qwen2.5-VL-72B [3] is included for reproduction. To comprehensively assess model capabilities on different languages, each image in our benchmark is paired with one English (EN) and one Chinese (CN) instruction. For EN instructions (GEdit-Bench-EN), both closed and open-source models are evaluated. For CN instructions (GEdit-Bench-CN), only those models which supports Chinese prompts, i.e., the close-source systems, are tested. During the evaluation process, we find that close-source image editing system such as GPT-4o may refuse certain instructions due to safety policies. To address this issue, we report two scores for two testing sets: (1) Intersection Subset — the subset of images whose results can be successfully returned from all the tested models, and (2) Full set — all the testing samples from GEdit-Bench. For the full set of results, we will calculate the average scores only for the cases where the models successfully generate and return the target image. For each evaluated model, instances where no result image is returned due to reasons such as safety concerns will be excluded from the averaging process.
@@
Fig. 6 demonstrates the groundbreaking capacity of Step1X-Edit, which outperforms open-source counterparts across 11 distinct evaluation axes. When compared to closed models, it surpasses Gemini2 Flash [14] and even beats GPT-4o [33] in axes such as style change and color alteration. As detailed in Tab. 2, Step1X-Edit significantly outperforms the existing open-source models like OmniGen [56] and has comparable results against the closed model like Gemini2 Flash and Doubao. Furthermore, as shown in Tab. 3, Step1X-Edit demonstrates consistent performance, even surpassing Gemini2 and doubao when handling the Chinese editing instructions in GEdit-Bench-CN benchmark. These results highlight the outstanding performance of our model across all dimensions with a unified architecture, eliminating the requirement for masks during the editing process. Fig. 7 and Fig. 8 provide illustrative examples for English instructions and Chinese instructions, respectively.

### v3

@@ 1 Introduction
Image editing with natural language instructions has become an increasingly important task in vision-language research. It offers intuitive interaction for end users while posing unique technical challenges: understanding nuanced semantics, precisely localizing regions to edit, and preserving image fidelity. While diffusion models [42, 5, 7, 11, 41] have dramatically improved image generation quality, the existing design by integrating text encoder, e.g., CLIP [44] and T5 [45], with diffusion transformer often struggles with following editing instruction to maintain alignment between input image and edit instruction, especially when edit instructions are subtle or compositional.
@@
Building on this dataset, we propose, Step1X-Edit, a unified image editing model that combines the strong semantic reasoning of Multimedia Large Language Model (MLLM), e.g., Qwen-VL [3], with a DiT-style diffusion architecture. The reference image and editing prompts can be processed by the MLLM to generate a target image latent condition which will be integrated with the diffusion model to obtain the output image. Our approach maintains a good balance between reference image reconstruction and editing prompt following. To train the model, we start from a text-to-image model to retain aesthetic quality and visual consistency, which can be easily replace by the existing text-to-image models like SD3 [11], FLUX[5, 7, 26], HiDream-I1 [18] and Flex [38]. To evaluate the existing editing models, we introduce a new benchmark named GEdit-Bench. By carefully collecting the images and editing prompts, GEdit-Bench ensures both real-world editing requirements and the diversity of the editing prompts. The experiments on GEdit-Bench validate that Step1X-Edit outperforms existing open-source baselines with a large-margin and approaches the performance of leading proprietary models, e.g., GPT-4o.
@@ 2.1 Controllable Image Generation and Edit
Diffusion models have become the dominant approach for high-fidelity image synthesis, offering strong capabilities in photorealism, structural consistency, and diversity. Beginning with DDPM [19] and DDIM [51], and further advanced by Latent Diffusion [47, 42], diffusion models operate in latent spaces for improved scalability. With the introduction of DiT architectures [41], diffusion models have made significant strides in generalization, image quality, and knowledge capacity, becoming the predominant architecture in modern image generation [1, 5, 7]. Based on the above text2image models, ControlNet [67], and T2I-Adapter [36] inject spatial or task-specific control into the generation process. BrushNet [23], PowerPaint [73], and FLUX-Fill [6] further improve inpainting quality and versatility. Despite these advances, diffusion models often rely on static prompts or fixed conditions and lack the capacity for multi-turn reasoning or flexible language alignment, limiting their application in open-ended editing scenarios.
@@
2.2 Instruction-based Image Editing Models
Instruction-based image editing models aim to bridge semantic instruction understanding and precise visual manipulation. Early approaches such as InstructEdit [55], InstructPix2Pix [8], MagicBrush [66], and BrushEdit [28] adopt modular pipelines where MLLMs generate prompts, spatial cues, or synthetic instruction-image pairs to guide diffusion-based editing.
Recent works move toward tighter integration between instruction and generation. SmartEdit [21], X2I [33], RPG [62], AnyEdit [64], and UltraEdit [71] enhance multimodal interaction and instruction fidelity through improved model architectures, task-aware routing, and fine-grained editing capabilities. Meanwhile, unified generation and editing frameworks such as OmniGen [61], ACE [16], ACE++[34], and Lumina-OmniLV[43] consolidate diverse visual tasks under a single architecture. Methods like Qwen2VL-Flux [31], DreamEngine [9] and MetaQueries [39] further explore efficient control integration and direct latent-level fusion between MLLMs and diffusion decoders. Moreover, Hidream-E1 [17] incorporates instructions and edited image descriptions as inputs, enabling more detailed edit information. Additionally, some researchers focus on more efficient methods to ensure editing performance while reducing training costs. For example, ICEdit [70] employs LoRA-MoE hybrid tuning and identifies better initial noise distributions, while SuperEdit [35] uses higher-quality training data and introduces contrastive supervision signals.
@@ 3.2 Our Method
Figure 5: DiT module details.
The extracted embeddings are then fed into a lightweight connector module, such as the token refiner [32, 24]. This module restructures the embeddings into a more compact multimodal feature representation, which is subsequently used as the multimodal embedding input to the downstream DiT network. Furthermore, we calculate the mean of valid embeddings from Qwen. This mean value is then projected through a linear layer, generating a global guidance vector. By doing so, the image editing network can leverage Qwen’s enhanced semantic comprehension capabilities, enabling more accurate and context-aware editing operations.
To effectively train the Token Refiner and enable rich cross-modal conditioning, we carefully design our feature aggregation strategy. Compared to approaches such as FLUX-Fill [6], which uses channel concatenation, and methods like SeedEdit [50], which introduce additional causal self-attention mechanisms, we follow OminiControl [53] and adopt token concatenation to better balance responsiveness to editing instructions with the preservation of fine-grained image details. During training, the reference image is encoded by a VAE encoder, and its latent features are linearly projected into reference image tokens. As illustrated in Fig. 5, the image tokens (highlighted in the green box) are concatenated with noise image tokens along the token length dimension to construct the final visual input.
Recent research has also explored alternative approaches to enhance cross-model understanding by leveraging MLLMs. For instance, Qwen2VL-FLUX [31] introduces a pioneering approach by replacing the traditional T5 [45] text encoder in DiT-based text-to-image models with MLLMs to enhance multi-modal understanding and generation. However, it still retains T5 [45] for text encoding, which limits its ability to perform comprehensive cross-modal reasoning. In contrast, approaches like DreamEngine [9] leverage Qwen to align image and text modalities, with their features serving as external conditional inputs for SD3.5 [1] in the image synthesis. This method establishes a shared representation space that facilitates a more coherent process of generation and understanding. Nevertheless, in DreamEngine, the challenge remains in fully capturing the fine-grained details of reference images using only MLLM features.
@@ 4.1 GEdit(Genuine Edit)-Bench
Table 2: Quantitative evaluation on GEdit-Bench-EN. All metrics are reported as higher-is-better ($\uparrow$). The Intersection subset reflects the subset of prompts where all methods return valid responses with a total of 434 instances; the Full set includes all the 606 instances. G_SC, G_PQ, and G_O refer to the metrics evaluated by GPT-4.1, while Q_SC, Q_PQ, and Q_O refer to the metrics evaluated by Qwen2.5-VL-72B.
@@
Table 3: Quantitative evaluation on GEdit-Bench-CN. All metrics are reported as higher-is-better ($\uparrow$). The Intersection subset reflects the subset of prompts where all methods return valid responses with a total of 422 instances; the Full set includes all the 606 instances. G_SC, G_PQ, and G_O refer to the metrics evaluated by GPT-4.1, while Q_SC, Q_PQ, and Q_O refer to the metrics evaluated by Qwen2.5-VL-72B.
@@ 4.2.1 Evaluation on GEdit-Bench
Based on the GEdit-Bench, we evaluated a diverse range of image editing algorithms, covering state-of-the-art open-source solutions such as Instruct-Pix2Pix [8], MagicBrush [66], AnyEdit [64], OmniGen [61], as well as proprietary algorithms like GPT-4.1 [37] [^1] The results are obtained based on ChatGPT APP in April 2025., doubao [50] [^2] The results are obtained based on doubao APP in April 2025., and Gemini2 Flash [15] [^3] The results are obtained in April 2025.. Following VIEScore [25], we adopt three metrics: SQ (Semantic Consistency), PQ (Perceptual Quality), and O (Overall Score). SQ assesses the degree to which the edited results conform to the given editing instruction, with a score ranging from 00 to $10$. PQ evaluates the naturalness of the image and the presence of artifacts, also using a scoring scale that ranges from 00 to $10$. The overall score is calculated based on these evaluations. To perform the automatic evaluation for VIEScore, we adopt the state-of-art MLLM model GPT-4o [^4] API access as of April 2025. Also, the evaluation based on the open-source model Qwen2.5-VL-72B [3] is included for reproduction. To comprehensively assess model capabilities on different languages, each image in our benchmark is paired with one English (EN) and one Chinese (CN) instruction. For EN instructions (GEdit-Bench-EN), both closed and open-source models are evaluated. For CN instructions (GEdit-Bench-CN), only those models which supports Chinese prompts, i.e., the close-source systems, are tested. During the evaluation process, we find that close-source image editing system such as GPT-4o may refuse certain instructions due to safety policies. To address this issue, we report two scores for two testing sets: (1) Intersection Subset — the subset of images whose results can be successfully returned from all the tested models, and (2) Full set — all the testing samples from GEdit-Bench. For the full set of results, we will calculate the average scores only for the cases where the models successfully generate and return the target image. For each evaluated model, instances where no result image is returned due to reasons such as safety concerns will be excluded from the averaging process.
@@
Fig. 7 demonstrates the groundbreaking capacity of Step1X-Edit, which outperforms open-source counterparts across 11 distinct evaluation axes. When compared to closed models, it surpasses Gemini2 Flash [15] and even beats GPT-4o [37] in axes such as style change and color alteration. As detailed in Tab. 2, Step1X-Edit significantly outperforms the existing open-source models like OmniGen [61] and has comparable results against the closed model like Gemini2 Flash and Doubao. Furthermore, as shown in Tab. 3, Step1X-Edit demonstrates consistent performance, even surpassing Gemini2 Flash and Doubao when handling the Chinese editing instructions in GEdit-Bench-CN benchmark. These results highlight the outstanding performance of our model across all dimensions with a unified architecture, eliminating the requirement for masks during the editing process. Fig. 8 and Fig. 9 provide illustrative examples for English instructions and Chinese instructions, respectively.